OpenStudy (anonymous):

cos(pie)/7cos(pie)/5-sin(pie)/7Sin(pie)/5 write the expression as sin, cos or tan of an angle

OpenStudy (jdoe0001):

wooow! so many PIES that's a really delicious expression

OpenStudy (mertsj):

\[\cos x \cos y-\sin x \sin y=\cos (x+y)\]

OpenStudy (mertsj):

Perhaps you could match your problem up with that identity.

OpenStudy (anonymous):

hey is the answer cos(2pie)/35

OpenStudy (mertsj):



OpenStudy (mertsj):

What is x?

OpenStudy (anonymous):

x is \[\pi / 7\]

OpenStudy (mertsj):

What is y?

OpenStudy (anonymous):

y is \[\pi/5\]

OpenStudy (mertsj):

What is \[\frac{\pi}{7}+\frac{\pi}{5}\]

OpenStudy (anonymous):

\[2\pi/35 \] \[\pi/12\]

OpenStudy (anonymous):

don't know if you have times it or add it

OpenStudy (mertsj):

\[\frac{\pi}{7}\times \frac{5}{5} +\frac{\pi}{5}\times \frac{7}{7}=\] \[\frac{5\pi}{35}+\frac{7\pi}{35}=\frac{12\pi}{35}\]

OpenStudy (anonymous):

oh ok thanks